OpenStudy (anonymous):

Which equation is parallel to the line y = −3x + 12 and passes through the point (−1, 6). y = 1/3x + 7 y = −3x + 3 y = 1/3 x + 3 y = −3x + 7

OpenStudy (apoorvk):

B. for parallel lines slope is same. so line-2 will be y=(1/3)x + k here k is the y intercept or the constant term now when you satisfy (-1,6) in the new lines equation (since it passes through it), you get k=3 so you get the equation B.!

OpenStudy (anonymous):

If the equation is parallel to: y = −3x + 12, then it would have the same slope as that equation. So, I'd start by identifying the slope of the give equation. The slope is the number in front of the x, so our slope is -3. From here, let's set up a new equation y=mx+b Since we know the slope of our new equation, we can put it in for m. y=-3x+b Now, using the point they gave us (-1,6), we can force the b out. The -1 is the x-coordinate and the 6 is the y-coordinate, so we can plug them in for x and y, and solve for b. (6)=-3(-1)+b Simplify 6=3+b Subtract 3 from both sides 3 = b So the final answer is y=-3x+3